# How to Build a Quantum Supercomputer: Scaling Challenges and Opportunities

### v1

@@ VI.2 Multi-QPU: Strongly disordered quantum spin glasses
This particular ACK method draws from tensor network (TN) approaches developed in the quantum physics and quantum chemistry communities White (1993); Meyer et al. (1990). Tensor networks represent quantum states in a compressed form, and can provide structure to characterize entanglement patterns. In the context of quantum circuits, a TN can be efficiently expressed as circuit of linear depth as was shown by Lin et. al Lin et al. (2021) for matrix product states (MPS), a type of TN widely used to study 1D quantum systems. Combining the structure of TNs with linear depth quantum circuits is the basis for this ACK method.
@@
Figure 37: Schematic of an adaptive circuit knitting method using tensor networks. In the inner loop, a variational optimizer finds circuit parameters for partitions of a quantum system (based on a tensor network) in parallel. In the outer loop, an adaptive procedure finds cuts which minimize entanglement between partitions. After the best cuts are found, observables are reconstructed via circuit knitting.
@@
Figure 37 provides a schematic for this adaptive circuit knitting method. It begins with a TN representation of a quantum state (in the figure, an MPS for a 2D spin lattice). The TN is partitioned into $N$ sub-networks, drawn here with $N=4$. In an inner loop, for each sub-network the variational procedure outlined by Lin et al. Lin et al. (2021) is followed and gates $U(\theta_{N})$ are optimized for an efficient circuit representation of each sub-network. Following this optimization, in the outer loop entanglement measures (e.g., von Neumann entropy) are computed and an entropy heatmap among the qubits is constructed. Although this heatmap is partial (we do not have entropy measures at the cuts between partitions), the information available is used to update cuts to locations with low entanglement. With new cuts on the following iterations, the blind spots are revealed. The adaptive outer loop exits when entanglement between partitions is minimized. The inner loop can be parallelized since each sub-network is independent, and both the inner and outer loop can be executed using classical HPC. Once optimal partitions have been found, a measured observable can be reconstructed from the sub-circuits via circuit knitting Piveteau and Sutter (2023). As we show below, cutting gates at locations of minimal entanglement can substantially lower the classical overhead of circuit knitting.
@@
Simulating quantum systems for materials science or quantum chemistry is one of the most promising applications for quantum computers. Spin-lattice systems are well-studied in materials science, and despite their simplicity can exhibit complex quantum phenomena that are difficult to simulate classically, e.g. the dynamical quantum phase transitions discussed in Sec VI.1. As a prototype system, we apply this ACK method to simulating the non-equilibrium dynamics of a strongly-disordered spin chain evolving under an Ising model with transverse and longitudinal fields given by the Hamiltonian
@@
where $\sigma^{z}$ and $\sigma^{x}$ are the Pauli $Z$ and $X$ matrices, respectively, $i$ indexes the lattice site, and $J$s, $g$s, and $h$s are real-valued parameters. We study strongly disordered systems (where parameters are varied at each lattice site) because they are important for understanding exotic states of matter, they can be difficult to study, and because they lead to many-body localization effects which we suspect could be exploited for more-efficient simulation.
Figure 38 provides a summary of the results for an ensemble of 32-qubit spin chains, each time-evolving under a Hamiltonian with different parameters chosen randomly from a uniform distribution on [$-1,1$] (excluding 0). For each spin chain, circuit optimization was carried out at eight timesteps throughout the dynamics. We partition each system into two sub-circuits and compare the overhead cost of circuit knitting for a naïve cut in the middle of the chain (a “load-balanced” choice) versus a cut recommended by the entropy heatmap from the adaptive algorithm. Figure 38(a) gives a schematic for a single instance on a smaller 20-qubit system. Figure 38(b) shows the distribution of overheads for reconstructing a magnetization observable via circuit knitting for the adaptive and load-balanced cuts. Both cuts achieve similar accuracy in the observable, but in most cases the adaptive cut results in a much lower overhead—the green distribution is distinctly shifted to the left. On a case-by-case basis, the median reduction in cost was 15$\times$, while the 75th and 95th percentiles were 59$\times$ and 450$\times$, respectively. While not shown in this figure, the gap between adaptive and load-balanced widens during the later timesteps, indicating that for longer simulations the benefits of adaptive circuit knitting will increase. As in Section VI.1, the cuStateVec simulator was employed to enable performant execution of the 32-qubit simulations.
Figure 38: (a) Example of a 20-qubit spin chain where the adaptive cut is chosen at the minimum entanglement entropy. (b) Histogram of sampling overheads resulting from adaptive and load-balanced baseline cuts for an ensemble of 32-qubit strongly-disordered spin chains.
@@
VII Toward heterogeneous quantum and probabilistic computing
Parallel to the development of NISQ devices and FTQCs architectures, an emerging trend in computing has been to build quantum-inspired accelerators for combinatorial optimization and sampling problems. A notable example is the notion of a probabilistic computer (p-computer) with probabilistic bits (p-bit) Camsari et al. (2017a, b); Aadit et al. (2022). It has been shown that networks of hardware p-bits natively represent a wide class of probabilistic algorithms typically implemented in software, with significant energy and performance benefits Sutton et al. (2020); Aadit et al. (2022); Niazi et al. (2024); Singh et al. (2024); Nikhar et al. (2024). From an HPC perspective, domain-specific probabilistic computers can accelerate hard optimization and sampling tasks by orders of magnitude. These accelerators can be integrated in a distributed fashion within heterogeneous hybrid quantum–classical hardware architectures, which we discuss further in Section VII.4.
@@ VII.1 Probabilistic computing with intrinsic higher-order interactions
There are two immediate generalizations possible: (a) p-bits can be extended to have multiple states. These Potts spins have also been implemented in hardware Whitehead et al. (2023) and shown to have better embedding than simple p-bits for certain optimization problems such as graph coloring. (b) The graph connectivity defined by $J_{ij}$ can be generalized to hypergraphs where the local field equation becomes (e.g., for $k=4$-local interactions)
@@ VII.2 Hardware implementation of p-computers
Physical implementation of p-computers includes a wide range of choices from noisy materials to analog and digital CMOS. State-of-the-art p-computers demonstrate nanodevice (magnetic tunnel junction, MTJ) based prototypes Borders et al. (2019); Singh et al. (2024), where the natural noise of the stochastic MTJ provides computational resources to solve a small-scale problem. This prototype has shown the promise of magnetic RAM technology, as a scalable pathway to build energy-efficient p-computers. Magnetic memory industry has achieved Gigabit densities of magnetic tunnel junctions embedded with CMOS transistors in monolithic integrated circuits Lee et al. (2019). Repurposing these MRAM chips so that their stable MTJs become unstable (low-barrier) could lead to dedicated probabilistic computers with tens of millions of integrated p-bits. Before an integrated p-computer using millions of stochastic magnetic MTJs, however, digital emulators of p-bits using powerful CMOS-based Field Programmable Gate Arrays (FPGA) have been used to investigate the architectural and algorithmic performance of p-computers at large-scale Sutton et al. (2020); Kaiser and Datta (2021); Aadit et al. (2022); Niazi et al. (2024); Nikhar et al. (2024); Singh et al. (2024). Even single FPGA-based implementations of p-computers have shown competitive performance against the state-of-the art Nikhar et al. (2024).
@@ VII.3 Scaling up p-computers: a distributed approach
Very often the sizes of single processing elements (PE), in the form of FPGAs/GPUs/TPUs, are not large enough to encode practical problem sizes containing thousands to millions of variables. To get around this problem, one solution is to design distributed architectures where a large problem is partitioned into smaller subgraphs housed in distinct PEs (Figure 40). Preliminary results show that as long as the communication links are faster than individual p-bit clocks, distributed probabilistic computers can create the “illusion” Radway et al. (2021) of a single PE that can house a much larger graph. Sampling rates over 1000 flips per nanosecond are feasible (in preparation), boding well for hard combinatorial optimization and sampling problems.
More generally, many workloads in probabilistic AI models, such as modern Energy-based Models (EBM) Liao et al. (2022); Niazi et al. (2024), involve several linear algebra operations, mainly as matrix multiplies. Thus, to scale up probabilistic architectures a heterogeneous approach involving traditional digital accelerators (e.g., GPU) and p-computers is desirable. The GPU can be used for a large part of the forward operation, such as computing embeddings, the gradients, and loss optimization while the p-computers (implemented for example on an FPGA) can essentially implement the stochastic neuron operation. Figure 41 shows the example of a heterogenous system architecture with multiple GPUs and FPGAs that can be used for scaling up to a very large number of p-bits (e.g. $>$1B) making it able to train and infer next-generation EBMs.
Two natural concerns arise. First, given the EBM consists of a lot of matrix operation, careful system-level profiling should be performed to assess if the time for sampling is the one to optimize by Amdahl’s law. Note that in the generalized case of a fully connected model, by scaling linearly the number of neurons the number of synapses scales quadratically so at each step of an exponentially long sampling process, linear operations (i.e. activations) are performed on neurons and quadratic operations (i.e. matrix multiplies) on synapses.
@@ VII.4 Quantum-assisted probabilistic computing with custom-design accelerators
VIII Discussions
VIII.1 Supply chain management toward utility-scale
@@
A utility-scale quantum computer should be characterized as a machine whose computational value surpasses its total cost Altepeter (2022). This cost encompasses not only manufacturing and operational expenses but also the amortized research and development investments. Achieving the utility scale is a crucial milestone in quantum computing, signifying the point at which these advanced systems become economically viable for practical applications. A key premise of our position is that leveraging the existing semiconductor supply chain can help amortize the cost of research and manufacturing. To further this amortization, it is also important to ensure reusing quantum modules and manufacturing technologies for as many quantum applications as possible, with significant computational demands.
@@
Analysis of the semiconductor supply chain raises two issues: 1) Is the supply chain competitive enough to reduce costs while avoiding dependence on a single source or foreign-controlled manufacturing? 2) What existing semiconductor research programs can be leveraged to reduce the research cost of a utility-scale quantum computer within next 10 years? From a manufacturing cost perspective, moving to semiconductor compatible fabrication processes is essential to drive down costs. A number of 300-mm advanced fabrication facilities are being built in the U.S. through public-private partnerships (e.g., the US CHIPS Act) which will reduce reliance on single-source/foreign manufacturers. Although this approach would require upgrading the tooling of the semiconductor foundries, the upgrades are a fraction of the cost of building a state-of-the-art 300-mm foundry. This in turn will drive down the cost of the various components needed to build a quantum supercomputer. Furthermore, existing research fabs at locations such as Applied Materials and NY CREATEs could act as a stopgap fabrication facility prior to full commercialization of quantum computing. Operating a quantum computer will likely be roughly comparable with operating GPU racks in the sense that quantum computers have sophisticated cooling requirements and can also be energy-intensive to operate. For example, today’s superconducting quantum computers require roughly 20 kW to operate, a majority of which is consumed by the dilution refrigerator and microwave electronics.
@@
A natural follow-up question for utility-scale computing is the estimated timeline for development. This will be addressed next.
VIII.2 Exponential scaling progress: A mirage or reality?
Figure 42: Schematic plot of the number of qubits from three experiments at UCSB and Google over time, which describes a Moore’s law growth in the number of qubits. After the quantum supremacy experiment in 2019 Arute et al. (2019), a target of one million qubits at the end of the decade was projected by industry leaders. However, the current pace of hardware progress suggests that goal might be postponed by several decades, assuming an optimistic scenario that none of the scaling challenges mentioned here slows or halts the progress. To arrive at utility-scale quantum computers in the 2030-2035 time frame, we need a major increase in the rate of progress over the next five years. Our thesis is that new fabrication and systems design as well as full-stack HPC integration are required to tackle this challenge. For estimates of future scaling, we suggest using the number of qubits that can be entangled in practice (which assumes sufficient qubit connectivity with multiqubit gates that are fast and accurate enough)Cao et al. (2023).
Since the quantum supremacy milestone Arute et al. (2019), there has been an expectation that scaling the number of qubits will follow a Moore’s law (exponential) growth over time from $\sim$50 qubits to a million qubits at the end of this decade. This goal would mark the arrival of a practical fault-tolerant quantum computer. We chart this progress in Figure 42, starting from 2014 when a repetition code experiment was performed at UCSB (nine qubits) Kelly et al. (2015), through the Google quantum supremacy on random circuits Arute et al. (2019) (53 qubits), on to the most recent surface-code error-correction experiment (105 qubits) Acharya et al. (2024). Care must be taken in plotting only the qubit quantity since this ignores other important qubit metrics such as quality, speed, and connectivity. Here, these data points represent qubit systems with some degree of consistency in these four key metrics, thus it is a fairly reasonable plot.
The trajectory of these three data points appears to follow an exponential improvement, but at a much lower slope than is needed to reach one million qubits by 2030. Larger numbers of qubits have been reported for superconducting quantum processors, but even for these more optimistic characterizations of functional qubits, the scaling falls short of the original industry expectation. Will it be possible to greatly increase the slope over the next five years, corresponding to a double exponential growth? We think this is unlikely because the added challenges for scaling beyond 1000 qubits—detailed in Section I.2—could hinder even staying on the current growth path.
As we have outlined in this position paper, one way to increase the rate of exponential growth is to identify all major technical challenges that are blocking progress and devise mitigation strategies. These include taking a radically different approach to qubit fabrication and developing a full-stack system integration with heterogeneous high-performance computing infrastructures.
We have also studied the detailed trade-off of physical and computational resources for scalable error-corrected quantum computers based on resource estimates of classically hard electronic structure calculations leveraging realistic performance characteristics for superconducting qubits. We showed for quantum simulations of FeMoco with chemical accuracy ϵ=italic-$\epsilon=$ 1.0 mHa, one requires hundreds of millions of physical qubits and a minimum of 2.5 years to run with state-of-the-art hardware quality (baseline hardware). The runtime can be reduced to approximately half a year and the qubit count to tens of millions if the hardware quality is significantly improved towards desired hardware. Improving the hardware quality from baseline to desired values results in a reduced runtime and qubit count of approximately a factor of five. Our sensitivity analysis shows that improvements in the gate-control errors yield the most significant impact, whereas improvements in SPAM errors and coherence enhancements are significantly less effective for achieving better performance in error suppression. These findings suggest that quantum gate fidelity improvements are much more important than SPAM or idling qubit error rates for scaling logical performance. Furthermore, we demonstrate the robustness of lattice surgeries spread among separate capacitively coupled QPU wafers and even separate QPUs distributed among multiple DRs. We conclude that distributed surface code architectures across multiple DRs can tolerate two-qubit errors on the order of $1\%$ arising from noisy optical interconnects between the DRs.

### v2

@@ VI.2 Multi-QPU: Strongly disordered quantum spin glasses
This particular ACK method draws from tensor network (TN) approaches developed in the quantum physics and quantum chemistry communities White (1993); Meyer et al. (1990). Tensor networks represent quantum states in a compressed form, and can provide structure to characterize entanglement patterns. In the context of quantum circuits, a TN can be efficiently expressed as circuit of linear depth as was shown by Lin et al.  Lin et al. (2021) for matrix product states (MPS), a type of TN widely used to study 1D quantum systems. Combining the structure of TNs with linear depth quantum circuits is the basis for this ACK method.
@@
Figure 40: Schematic of an adaptive circuit knitting method using tensor networks. In the inner loop, a variational optimizer finds in parallel circuit parameters for partitions of a quantum system that are induced by an initial TN ansatz. In the outer loop, an adaptive procedure finds cuts which minimize entanglement between partitions. After the best cuts are found, quantum observables are reconstructed via circuit knitting. The procedure is repeated until a desired accuracy of some (global) observables is achieved or the performance of circuit knitting is plateaued.
@@
Figure 40 shows a schematic for this adaptive circuit knitting method. It begins with a TN representation of a quantum state (in the figure, an MPS for a 2D spin lattice). The TN is partitioned into $N$ sub-networks, drawn here with $N=4$. In an inner loop, for each sub-network the variational procedure outlined by Lin et al. Lin et al. (2021) is followed and gates $U(\theta_{N})$ are optimized for an efficient circuit representation of each sub-network. Following this optimization, in the outer loop entanglement measures (e.g., von Neumann entropy) are computed and an entropy heatmap among the qubits is constructed. Although this heatmap is partial (we do not have entropy measures at the cuts between partitions), the information available is used to update cuts to locations with low entanglement. With new cuts on the following iterations, the blind spots are revealed. The adaptive outer loop exits when entanglement between partitions is minimized. The inner loop can be parallelized since each sub-network is independent, and both the inner and outer loop can be executed using classical HPC. Once optimal partitions have been found, a measured observable can be reconstructed from the sub-circuits via circuit knitting Piveteau and Sutter (2023). As we show below, cutting gates at locations of minimal entanglement can substantially lower the classical overhead of circuit knitting.
@@
Simulating quantum systems for materials science or quantum chemistry is one of the most promising applications for quantum computers. Spin-lattice systems are well-studied in materials science, and despite their simplicity can exhibit complex quantum phenomena that are difficult to simulate classically, for example, the dynamical quantum phase transitions discussed in Section VI.1. As a prototype, we apply this ACK method to simulating the non-equilibrium dynamics of a strongly disordered spin chain evolving under an Ising model with transverse and longitudinal fields given by the Hamiltonian
@@
where $\sigma^{z}$ and $\sigma^{x}$ are the Pauli $Z$ and $X$ matrices, respectively, $i$ indexes the lattice site, and $J$s, $g$s, and $h$s are real-valued parameters. We study strongly disordered systems (where parameters are varied at each lattice site) because they are important for understanding exotic states of matter, they can be difficult to study, and because they lead to many-body localization effects which we hope to exploit for more efficient simulation.
Figure 41 provides a summary of the results for an ensemble of 32-qubit spin chains, each time-evolving under a Hamiltonian with different sets of parameters $J_{i,i+1}$, $g_{i}$, $h_{i}$ chosen randomly from a uniform distribution on [$-1,1$] (excluding 0). For each spin chain, circuit optimization was carried out at eight time steps throughout the dynamics. We partition each system into two sub-circuits and compare the overhead cost of circuit knitting for a naïve cut in the middle of the chain (a “load-balanced” choice) versus a cut recommended by the entropy heatmap from the adaptive algorithm. Figure 41(a) gives a schematic for a single instance on a smaller 20-qubit system. Figure 41(b) shows the distribution of overheads for reconstructing a magnetization observable via circuit knitting for the adaptive and load-balanced cuts. Both cuts achieve similar accuracy in the observable, but in most cases the adaptive cut results in a much lower overhead—the green distribution is distinctly shifted to the left. On a case-by-case basis, the median reduction in cost was 15$\times$, while the 75th and 95th percentiles were 59$\times$ and 450$\times$, respectively. While not shown in this figure, the gap between adaptive and load-balanced widens during the later time steps, indicating that for longer simulations the benefits of adaptive circuit knitting will increase. As in Section VI.1, the cuStateVec simulator was employed to enable performant execution of the 32-qubit simulations.
Figure 41: (a) Example of a spin chain where the adaptive cut is chosen at the minimum entanglement entropy. (b) Histogram of sampling overheads resulting from adaptive and load-balanced baseline cuts for an ensemble of 32-qubit strongly disordered spin chains.
@@
VII Toward quantum-inspired and quantum-assisted probabilistic computing
The power of randomness in computational algorithms has long been recognized by computer scientists. In particular, Monte Carlo algorithms, first popularized by Ulam and von Neumann and later developed by Metropolis and Teller, have been widely recognized as one of the top ten algorithms of the 20th century. Today, a considerable amount of workloads in combinatorial optimization, computational science, and AI are intrinsically probabilistic. While such workloads could in principle be accelerated by QPUs (e.g., quantum annealing for combinatorial optimization), shorter-term deployment of industry-scale probabilistic computers could be an efficient way to develop software stacks, know-how, and product market fit that would be quantum-ready and thus can be seamlessly integrated by future quantum technologies. Moreover, it is not clear why we should not capitalize on the probabilistic nature of classical many-body systems to significantly enhance our computational models and frameworks. This allows us to develop novel architectures leveraging intrinsically noisy physical systems, thus avoiding emulation of such probabilistic behavior on deterministic Turing machines, which can be expensive. For example, constructing an individual high-quality tunable random number generator could require up to 15,000 transistors per node Singh et al. (2024). More importantly, emulating many-body emergent probabilistic phenomena could be exponentially hard with the sequential Markov Chain Monte Carlo techniques, due to their long mixing times to arrive at non-trivial steady states.
Indeed, Feynman, who pioneered the notion of a controllable quantum computer for simulating other quantum systems (in his 1981 talk “Simulating Physics with Computers” Feynman (1982)), also imagined a probabilistic computer with the same idea: “The other way to simulate a probabilistic Nature $\ldots$ is by a computer $C$, which itself is probabilistic”. Motivated by this observation, a growing community has recently been developing a full-stack probabilistic computing approach Chowdhury et al. (2023a). Such probabilistic computers can be envisioned to take a great deal of inspiration from physics, where nature employs microscopic stochastic local dynamics, quasi-sparse graph connectivity, strong heterogeneity, non-equilibrium dynamics, and asynchronous updates (without global clocks) in probabilistic networks, see Fig 42. These networks evolve in a massively parallel manner, leading to emergent many-body phenomena at mesoscopic scales that are computationally rich and hard to simulate with conventional computing paradigms. Similar to GPUs that are designed to accelerate ML problems and QPUs specialized to simulate and model many-body quantum systems, many-body interacting probabilistic bits (p-bits) could be engineered to act as native processors for a wide range of probabilistic applications including Monte Carlo algorithms, Bayesian learning and inference Aifer et al. (2024a), training and inference in energy-based models Coles et al. (2023); Huembeli et al. (2022), combinatorial optimization, modeling complex systems, efficient linear algebra Aifer et al. (2024b), accelerating certain quantum simulations, and complex reasoning in generative AI Esencan et al. (2024).
Figure 42: Physical analogy between asynchronously interacting (a) bodies and (b) probabilistic bits: both systems are asynchronous, locally interacting with sparse connectivity and massively parallel.
While p-bits cannot substitute for qubits due to the well-known “sign problem” Chowdhury et al. (2023b), the boundary between probabilistic and quantum approaches is often blurry and continuously evolving. Specialized p-bit hardware and advanced sampling techniques can help push the boundaries of classical simulability, making these efforts highly worthwhile. It has been shown that networks of hardware p-bits natively represent a wide class of probabilistic algorithms typically implemented in software, with significant energy and performance benefits Sutton et al. (2020); Aadit et al. (2022); Niazi et al. (2024); Singh et al. (2024); Nikhar et al. (2024). From an HPC perspective, domain-specific probabilistic computers can accelerate hard optimization and sampling tasks by orders of magnitude. These accelerators can be integrated in a distributed fashion within high-performance heterogeneous hybrid quantum–classical hardware architectures, as we discuss in Section VII.4.
@@ VII.1 Probabilistic computing with intrinsic higher-order interactions
There are two immediate generalizations that can be made. One is that p-bits can be extended to have multiple states. These Potts spins have also been implemented in hardware Whitehead et al. (2023) and shown to have better embedding than simple p-bits for certain optimization problems such as graph coloring. The other is that the graph connectivity defined by $J_{ij}$ can be generalized to hypergraphs where the local field equation becomes (e.g., for $k=4$-local interactions)
@@ VII.2 Hardware implementation of p-computers
Physical implementation of p-computers includes a wide range of choices from noisy materials to analog and digital CMOS. State-of-the-art p-computers demonstrate nanodevice (magnetic tunnel junction, MTJ) based prototypes Borders et al. (2019); Singh et al. (2024), where the natural noise of the stochastic MTJ provides computational resources to solve a small-scale problem. This prototype has shown the promise of magnetic RAM technology, as a scalable pathway to build energy-efficient p-computers. Magnetic memory industry has achieved Gigabit densities of magnetic tunnel junctions embedded with CMOS transistors in monolithic integrated circuits Lee et al. (2019). Repurposing these MRAM chips so that their stable MTJs become unstable (low-barrier) could lead to dedicated probabilistic computers with tens of millions of integrated p-bits. Before an integrated p-computer using millions of stochastic magnetic MTJs, however, digital emulators of p-bits using powerful CMOS-based Field Programmable Gate Arrays (FPGA) have been used to investigate the architectural and algorithmic performance of p-computers at large-scale Sutton et al. (2020); Kaiser and Datta (2021); Aadit et al. (2022); Niazi et al. (2024); Nikhar et al. (2024); Singh et al. (2024). Even single FPGA-based implementations of p-computers have shown competitive performance against the state of the art Nikhar et al. (2024).
@@ VII.3 Scaling up p-computers: A distributed approach
Single processing elements (PE), such as FPGAs/GPUs/TPUs, are often not large enough to encode practical problem instances with thousands to millions of variables. To get around this problem, one solution is to design distributed architectures where a large problem is partitioned into smaller subgraphs housed in distinct PEs (Figure 44). Preliminary results show that as long as the communication links are faster than individual p-bit clocks, distributed probabilistic computers can create the “illusion” Radway et al. (2021) of a single PE that can house a much larger graph. Sampling rates over 1000 flips per nanosecond are feasible (manuscript in preparation), which bodes well for hard combinatorial optimization and sampling problems.
More generally, many workloads in probabilistic AI models, such as modern energy-based models (EBM) Liao et al. (2022); Niazi et al. (2024), involve several linear algebra operations, mainly as matrix multiplies. Thus, to scale up probabilistic architectures a heterogeneous approach involving traditional digital accelerators (e.g., GPU) and p-computers is desirable. The GPU can be used for a large part of the forward operation, such as computing embeddings, the gradients, and loss optimization while the p-computers (implemented for example on an FPGA) can essentially implement the stochastic neuron operation. Figure 45 illustrates a heterogenous system architecture with multiple GPUs and FPGAs that can be used for scaling up to a large number of p-bits (e.g., $>$1B) enabling it to train and infer next-generation EBMs.
Two natural concerns arise. First, given the EBM consists of a lot of matrix operation, careful system-level profiling should be performed to assess if the time for sampling is the one to optimize by Amdahl’s law. Note that in the generalized case of a fully connected model, by scaling linearly the number of neurons the number of synapses scales quadratically so at each step of an exponentially long sampling process, linear operations (i.e., activations) are performed on neurons and quadratic operations (i.e., matrix multiplies) on synapses.
@@ VII.4 Quantum-assisted probabilistic computing with custom-design accelerators
VIII Toward a cost-effective quantum computing technology
A utility-scale quantum computer should be characterized as a machine whose computational value surpasses its total cost Altepeter (2022). Achieving the utility scale is a crucial milestone in quantum computing, signifying the point at which these advanced systems become economically viable for practical applications. It is widely expected that initial quantum computers will be exorbitantly expensive, far beyond utility scale. Therefore we must analyze the rate at which the costs decrease for a quantum computer to cover more utility scale problems. We first describe the systems engineering considerations for reducing the cost of a quantum supercomputer. Then we discuss how the supply chain for the quantum supercomputer will reduce the costs of R&D and manufacturing over time.
VIII.1 Systems engineering
Based on the resource estimates from FeMoco, as a widely studied candidate for beyond-classical quantum simulation (see Section IV, high-level estimates suggest that a utility-scale quantum computer would require 10-50 quantum accelerators with each accelerator requiring between 100 thousand to 1 million qubits. These are rough estimates, and there is a fair amount of uncertainty due to the qubit quality, algorithmic improvements, technology that will be used to network the quantum accelerators, etc. An optimistic estimated cost to build each accelerator would be on the order of 10 million USD, hence one can imagine that a quantum supercomputer will cost on the order of 100 million to 1 billion USD.
Today, the costs of superconducting qubits are dominated by coaxial wiring, dilution refrigerators (due to size and power of circulators), and quality of the qubits. For example, a superconducting qubit quantum computer with 150 qubits roughly requires 4 million USD in just wiring, and 1 million USD for the dilution refrigerator. Brute force scaling to a 100k qubit system would be cost prohibitive. One key aspect of our approach is that the coaxial wiring can be eliminated through a combination of flex-wiring and the wafer scale wiring wafer. The quality of the qubits is also a key factor because the number of logical qubits that can be yielded is primarily driven by the quality of the qubits.
Additional aspects of our system can drastically reduce costs: cryogenic control leveraging off the shelf CMOS processes can reduce the cost per RF channel. Also, by leveraging existing HPC infrastructures, one can offload costly classical computation such as quantum error correction. For example, Table 8 shows that the computational power of the Nvidia DGX Quantum grows significantly in one generation.
VIII.2 Supply chain management
@@
The cost of a quantum computer encompasses not only manufacturing and operational expenses but also the amortized research and development investments. premise of our position is that leveraging the existing semiconductor supply chain can help amortize the cost of research and manufacturing. To further this amortization, it is also important to ensure reusing quantum modules and manufacturing technologies for as many quantum applications as possible, with significant computational demands.
@@
Analysis of the semiconductor supply chain raises two issues: 1) Is the supply chain competitive enough to reduce costs while avoiding dependence on a single source or foreign-controlled manufacturing? 2) What existing semiconductor research programs can be leveraged to reduce the research cost of a utility-scale quantum computer within next 10 years? From a manufacturing cost perspective, moving to semiconductor compatible fabrication processes is essential to reduce costs. A number of 300-mm advanced fabrication facilities are being built in the U.S. through public-private partnerships (e.g., the US CHIPS Act) which will reduce reliance on single-source/foreign manufacturers. Although this approach would require upgrading the tooling of the semiconductor foundries, the upgrades are a fraction of the cost of building a state-of-the-art 300-mm foundry. This in turn will reduce the cost of the various components needed to build a quantum supercomputer. Much of the reduction of costs of semiconductor products is a result of intense competition, therefore quantum computers that leverages the competitive semiconductor supply chain we will also see a reduction in costs over time. Furthermore, existing research fabs at locations such as Applied Materials and NY CREATEs could act as a stopgap fabrication facility prior to full commercialization of quantum computing. Operating a quantum computer can be compared with operating GPU racks in terms of sophisticated cooling requirements and energy-intensive operation. For example, today’s superconducting quantum computers need roughly 20 kW to operate, most of which is consumed by the dilution refrigerator and microwave electronics. Therefore, it is also imperative to build supply chains to lower the operating burden of quantum computers.
@@
IX Concluding remarks
The ultimate goal of developing a utility-scale quantum computer is to deliver valuable computations that justify the device’s cost. Achieving this objective for any architecture involves demonstrating several non-trivial criteria: (1) the existence of a set of computations that produce measurable value when executed; (2) the cost of a quantum computer capable of performing these computations is lower than the value they generate; and (3) no alternative methods – whether classical, analog, quantum, experimental, etc – can execute these computations more cost-effectively. Evaluating any proposal for a utility-scale quantum computer, including the vision we present here, is thus a challenging task involving answering questions spanning domain-specific use cases, device architecture and integration, and the full range of computational methods. This work focuses primarily on the second criterion, providing an initial holistic understanding of the size, scale, and cost of a quantum supercomputer and identifying key challenges.